OpenStudy (anonymous):

Find all polar coordinates of point P = (9, 75°).

OpenStudy (anonymous):

For cartesian to polar: r2=x2+y2 θ=tan−1(yx)

OpenStudy (anonymous):

I know that what to do (I think) I just need to make sure it's right.

OpenStudy (misssunshinexxoxo):

(9,75-360) is the correct answer

OpenStudy (misssunshinexxoxo):

N covers an integer

OpenStudy (anonymous):

so would it be (9, 75 -360n)? or just what you wrote? Also can you explain why?

OpenStudy (anonymous):

@misssunshinexxoxo

OpenStudy (misssunshinexxoxo):

(9,75+360n) where n is a integer covering it

OpenStudy (anonymous):

covering it?

OpenStudy (misssunshinexxoxo):

n is is a negative integer

OpenStudy (anonymous):

ok